The waiting time in line at an ice cream shop has a uniform distribution between 0 and 9 minutes. What is the 80th percentile of this distribution? (Recall: The 80th percentile divides the distribution into 2 parts so that 80% of area is to the left of 80th percentile) _______ minutes Answer: (Round answer to two decimal places.)
Answer #1

Answer: 7.2 minutes

Let k =80th percentile

X ~U(0,9)

So the probability density function

f(x) = 1/(9-0) = 1/9 for 0 <x < 9

k= 0.80*9

k = 7.20

Hence the 80 th percentile is 7.20 minutes.

Please do the comment for any doubt or clarification. Please upvote if this helps you out. Thank You!